OpenStudy (anonymous):

A 20% acid solution is mixed with a 70% acid solution to get 50 liters of a 40% solution. Which of the following expressions is equal to 4(50)? 0.2x + 0.7y 2x + 7y x + y

OpenStudy (anonymous):

@seb.cal

OpenStudy (anonymous):

@Callisto

OpenStudy (just_one_last_goodbye):

x * 20 + y * 70 = 40 * 50 x + y = 50 20x + 70y = 2000 2x + 7y = 200 x = 50 - y 2 * (50 - y) + 7y = 200 100 - 2y + 7y = 200 5y =100 y = 20 x = 50 - y x = 50 - 20 x = 30 30 liters of 20% 20 liters of 70%

OpenStudy (anonymous):

so what is it?